Minutes of Smathers
Middle Manager's Meeting
February 20, 1997

Present: Suzanne Brown, Tom Cetwinski, Gary Cornwell, Robena Cornwell, Stephanie Haas, Carolyn Henderson, Yael Herbsman, Dot Hope, Martha Hruska, John Ingram, Dolores Jenkins, Mari Marsh, Tom Minton, Richard Phillips, Colleen Seale, Carol Turner

General Announcements
  1. Dolores Jenkins touts the site, "Luddites On-Line" - http://www.luddites.com - as the only place in cyberspace devoted exclusively to luddites, technophobes and other refugees from the Information Revolution.
General Wiring
  1. MSL wiring should be complete. There are reports of some work taking place in room 206, but Bill was informed that they would be finished by this Tuesday and out by this Wednesday.
  2. Router work continues - CISCO should be tiring of 6:00 AM repair visits. Next time they will bring a new router if the problem returns. It is functional at the moment.
  3. The Health Center has been unable to access their CD's. Software patches have been installed and seem to work.
  4. No word on the campus fiber expansion.
  5. AFA to Music wiring - half of the transceivers for the job have been received.
  6. Campus computer reorganization structure - no word.
  7. News in the industry - Apple continues to slide.
  8. Biosis disk is here, but due to an indemnity clause and copyright problems, we are unable to use it. Biosis is working on a third party agreement with language they are comfortable in using. Expect this sort of problem to multiply as disks get used in unapproved ways - the companies will become irate (much like the recent Compustat tape debacle).
  9. A circuit breaker on the 2nd floor of Library West was tripped last night at 6:45 PM. PPD did not show, Murdoch has been notified. Bill urges that someone responsible who knows how to deal with such problems should have access to the power rooms where the circuit breakers are located.
  1. Cow herd of 80 boxes has arrived. Should be ready for distribution in the next day or so. These have already been checked for possible video problems. We will try to distribute them quickly - most are to go to staff. The next 40 to be ordered are in the pipeline and should be arriving sometime next week.

    Back to the Systems Homepage

    Last updated February 18, 2000
    by Debra Harris